Mini-Prj 2. Extraction of Thevenin and Norton Equivalent Circuits by LTspice Part A. Wheatstone Bridge Circuit with a VoltageII. Extraction of Thevenin Equivalent Circuit for this Wheatstone Bridge Circuit RTh VTh RL (c) Determining Thevenin equivaleR5 R1 R2 Ie Vs 0 R3 R4 (S2) Submit Answer Tries 0/3 (e) With the extracted VTh and RTh and using Thevenin equivalent circuit,III. Extraction of Norton Equivalent Circuit for this Wheatstone Bridge Circuit IN RN RL (g) Determining Norton equivalent cuMini-Prj 2. Extraction of Thevenin and Norton Equivalent Circuits by LTspice Part B. Wheatstone Bridge Circuit with a CurrentII. Extraction of Thevenin Equivalent Circuit for this Wheatstone Bridge Circuit RTh VTh RL (c) Determining Thevenin equivaleR5 R1 R2 Is Ie R3 R4 0 Submit Answer Tries 0/3 (e) With the extracted VTh and RTh and using Thevenin equivalent circuit, reca

Mini-Prj 2. Extraction of Thevenin and Norton Equivalent Circuits by LTspice Part A. Wheatstone Bridge Circuit with a Voltage Source Vs R5 R1 R2 Vs RL R3 R4 For the circuit as shown below, given that R1= 9 Ω, R2= 17 Ω, R3= 9 Ω' R,-18 Ω, R5= 19 Ω , RL= 2 Ω ,V,-74 V I. Wheatstone Bridge Circuit Analysis (a) Determining the load voltage Vi-Vab for the Wheatstone bridge circuit with LTspice Submit Answer Tries 0/3 (b) Determining the load current Iu following from a to b for the Wheatstone bridge circuit with LTspice. Submit Answer Tries 0/3
II. Extraction of Thevenin Equivalent Circuit for this Wheatstone Bridge Circuit RTh VTh RL (c) Determining Thevenin equivalent voltage VTh for this Wheatstone bridge circuit, where VTh is extracted from the open circuit voltage at the terminals a, b without the load, RL Submit Answer Tries o/3 (d) As shown below, extracting Thevenin equivalent resistance RTh for this Wheatstone bridge circuit by setting Vs-0 V and adding an external current source with Le-1.2 A. Thus, Thevenin equivalent resistance can be evaluated via RThV (Question: why not use Vab instead?)
R5 R1 R2 Ie Vs 0 R3 R4 (S2) Submit Answer Tries 0/3 (e) With the extracted VTh and RTh and using Thevenin equivalent circuit, recalculate the load voltage Vu-VLTh-Vab for the Wheatstone bridge circuit with LTspice VLTh Submit Answer Tries 0/3 ) With the extracted VTh and RTh and using Thevenin equivalent circuit, recalculate the load current IL-ILTh for the Wheatstone bridge circuit with LTspice. ILM" Submit Answer Tries 0/3
III. Extraction of Norton Equivalent Circuit for this Wheatstone Bridge Circuit IN RN RL (g) Determining Norton equivalent current IN for this Wheatstone bridge circuit, where IN is extracted from the source transformation IN VTh/RTh IN- Submit Answer Tries 0/3 (h) Determining Norton equivalent resistance RN, where RN is extracted from the source transformation RN-RTh RN Submit Answer Tries 0/3 (i) with the extracted IN and RN and using Norton equivalent circuit, recalculate the load voltage VL-VLN=Vab for the Wheatstone bridge circuit with LTSpice. VLN Submit Answer Tries 0/3 ) With the extracted IN and RN and using Norton equivalent circuit, recalculate the load current IL ILN for the Wheatstone bridge circuit with LTspice. Submit Answer Tries 0/3 ILN
Mini-Prj 2. Extraction of Thevenin and Norton Equivalent Circuits by LTspice Part B. Wheatstone Bridge Circuit with a Current Source Is R5 R1 R2 Is RL R3 R4 For the circuit as shown below, given that R1-22 Ω, R2-15 Ω, R3-28 Ω, R4-9 Ω, R5-29 Ω , R.-16 Ω, 1,-6 A. I. Wheatstone Bridge Circuit Analysis (a) Determining the load voltage VL-Vab for the Wheatstone bridge circuit with LTspice. Submit Answer Tries 0/4 (b) Determining the load current IL following from a to b for the Wheatstone bridge circuit with LTspice Submit Answer Tries 0/4
II. Extraction of Thevenin Equivalent Circuit for this Wheatstone Bridge Circuit RTh VTh RL (c) Determining Thevenin equivalent voltage VTh for this Wheatstone bridge circuit, where VTh is extracted from the open circuit voltage at the terminals a, b without the load, RL Submit Answer Tries 0/3 d) As shown below, extracting Thevenin equivalent resistance RTh or this wheatstone bridge circuit by setting 1s=0 A and adding an external current source with e=1.2 A. Thus, Thevenin equivalent resistan can be evaluated via RTh Va/e (Question: why not use Vab instead?)
R5 R1 R2 Is Ie R3 R4 0 Submit Answer Tries 0/3 (e) With the extracted VTh and RTh and using Thevenin equivalent circuit, recalculate the load voltage V VLTh-Vab for the Wheatstone bridge circuit with LTspice. Submit Answer Tries 0/3 (f) With the extracted VTh and RTh and using Thevenin equivalent circuit, recalculate the load current ILILTh following from left to right for the Wheatstone bridge circuit with LTspice ILTh Submit Answer Tries 0/3
III. Extraction of Norton Equivalent Circuit for this Wheatstone Bridge Circuit IN RN RL (g) Determining Norton equivalent current IN for this wheatstone bridge circuit, where IN įs extracted from the source transformation 1N= Vrh/RT- 1 (A) Submit Answer Tries 0/3 (h) Determining Norton equivalent resistance RN, where RN is extracted from the source transformation RN RN- RTh Submit Answer Tries 0/3 ) With the extracted IN and RN and using Norton equivalent circuit, recalculate the load voltage VL-VLN-Vab for the Wheatstone bridge circuit with LTspice VUN Submit Answer Tries 0/3 ) With the extracted IN and Rv and using Norton equivalent circuit, recalculate the load current IL ILN following from left to right for the Wheatstone bridge circuit with LTspice. ILN Submit Answer Tries 0/3
Answer #1

I. I have simulated the given wheat stone bridge using LTspice and results are shown below

LTspice XVII Draft1 Eile View Plot Settings Simulation Iools Window Help Egi Adi刍ㄧㄥ女叩之 3 D-纱屮り EmE] Aa op Draftt R5 19 R2 17LTspice XVII - Draft Eile View Plot Settings Simulation Iools Window Help ] Τ 茅49 Q Q皋賒ig E 義더 R龉生刍 / 回ぐ 3 D-epY Of E Aa.op

Waveforms for V(a) and V(b) can be shown on the same graph,but I have shown them on different graphs in this first step.

From simulations we have, V(a)=14.358 V and V(b)=14.411 V = > therefore V(ab)=V(a)-V(b)=14.358-14.411 = -0.052 V

Therefore, V(ab) = VL= -0.052 V.

LTspice XVII - Draftl File View Plot Settings Simulation Iools Window Help Draftt Draft R5 19 R2 Vs RL 74 V R3 R4 18 .tran 1

From simulation graph .Load current I(RL)=26.688 mA

Therefore, I(RL)=IL=0.0268 A

2. Now finding Vth by removing RL , we have

LTspice XVII - Draftl File View Plot Settings Simulation Iools Window Help Drat1 Draft1 Draft1 R5 19 R1, R2 9 Vs 74 V R4 18 R

Va(black line in graph)=14.24 V, Vb(Pink line in graph)=14.64 V

Therefore , Vth= Va - Vb = 14.24 - 14.64 = -0.4 V

Therefore Vth= -0.4 V

For Rth, current source=1.2 A is applied between a and b as shown below

LTspice XVII Draft1 Eile View Plot Settings Simulation Iools Window Help LDrDraft Draft1 R5 19 R1 R2 17 9 Vs 74 V 1.2 R4 18 RTherefore Vba=Vb - Va = 25 - 9 = 16 V , Current source (I1)=1.2 A

Rth=Vba/I1 = 16/1.2 = 13.33 ohm

Therefore , Rth=13.33 ohm

Therefore Thevenin equivalent circuit we have ,

LTspice XVII - Draft2 File View Plot Settings Simulation Tools Window Help Draft22 raft2 Rth 13.33 Vth RL 2 0.4 .tran 1 V(a) Vb=0 V(gnd), Va(black line)= -52.2 mV= -0.052 V , therefore Va - Vb= Vab=VLth= -0.052 V ,Hence same as above calculated.

IL= 26.2 mA(pink line in graph)

Hence ILth= 0.0262 A (nearly same as above calculated)

Now Norton equivalent circuit we have , In=Vth/Rth= -0.4/13.33 = -0.03 A , Rn=Rth= 13.33 ohm

LTspice XVII - Draft3 File View Plot Settings Simulation Tools Window Help DrafDraDraf2Draft2DratDraft? Draft3 In RL 13.33 2

Therefore from graph we have VLn=Va - Vb= -52.2mV -0 = -52.2 mV , therefore VLn= - 0.052 V

ILn=0.0262 A

As per HomeworkLib policy I have solved first question as you have posted multiple question. feel free to ask if you have any query.Thank you